OpenStudy (anonymous):

-x+4<2(x-8) (A INQUALITY PROBLEM)

OpenStudy (anonymous):

\[-x + 4 \lt 2x - 16\]\[-x + 4 - 4 \lt 2x - 16 -4\]\[-x \lt 2x - 20\]\[-x - 2x \lt 2x -2x -20\]\[-3x \lt -20\]\[\frac{-3x}{-3} \lt \frac{-20}{-3}\]\[x \gt \frac{20}{3}\]